OpenStudy (joshoyen):

Solve each system of equations. 1. y = -3x - 21 y = 5x + 19

pooja195 (pooja195):

y = -3x - 21 <---e1 y = 5x + 19 <---e2 Use substitution :-) Plug e1 into e2 since the value of y is given -3x-21=5x+19 solve for x

OpenStudy (joshoyen):

ohhh okay thanks!

pooja195 (pooja195):

Did you solve?

OpenStudy (joshoyen):

yep x = -5, y=-6

OpenStudy (joshoyen):

@pooja195 i'm actually stuck on this other problme, could you hlep me, it's 5x-4y=-19 x+5y=-27

pooja195 (pooja195):

Im more fimiliar with substituion x+5y=-27 subtract 5y from both sides now you have x=-5y-27 now you have 2 equations 5x-4y=-19 <---e1 x=-5y-27 <---e2 Plug e2 into e1 since the value of x is given 5(-5y-27)-4y=-19 and solve for y from there :)

OpenStudy (joshoyen):

thanks!

pooja195 (pooja195):

yw let me kno wt u get

OpenStudy (joshoyen):

x= -7 y= -4

pooja195 (pooja195):

Good job :-)